# Flipping Out: Role of Arginine in Hydrophobic Polymer Collapse

### v1

@@ Abstract
Maintaining native protein structures in biological formulations poses a challenge, and is commonly addressed by strategic additive incorporation.1, 2, 3 Arginine stands out as a frequently employed additive in such formulations, spanning both therapeutic proteins4 and vaccines.5, 6 Arginine has been widely used as an aggregation suppressor, an agent for protein refolding, a cryoprotectant during lyophilization, and in protein purification. 7, 8, 9 Once hailed as a universal stabilizer, emerging studies paint a foggier picture of the effects of arginine. In some settings, the presence of arginine has accelerated the aggregation 10, 11, 12, denaturation13, 14, 15, and activation16 of certain proteins and viruses. Additional studies have found that arginine increases or decreases protein melting temperature depending on concentration 17, 18. Hence, the existing literature on arginine reveals a lack of a cohesive understanding of its multi-faceted effects on protein stability.
@@
We simulated a hydrophobic polymer in arginine solutions at different concentrations (Table S2). Replica exchange umbrella sampling60 (REUS) simulations were performed using GROMACS 2021.461, 62 with the PLUMED 2.8.0 63, 64 patch applied. The hydrophobic polymer was modeled as a linear coarse-grained chain with 26 monomers, where each monomer represents a $\mathrm{CH_{2}}$ unit with Lennard-Jones parameters $\sigma=0.373$ $\mathrm{nm}$ and $\epsilon=0.5856$ $\mathrm{kJ/mol}$38. Chloride (${}^{-}$) counterions were added to achieve charge neutrality. The TIP4P/200565 model was used to describe water, and the CHARM22 force field was used for arginine and ${}^{-}$ 66. Lorentz-Berthelot mixing rules67 were used to calculate non-bonded interactions between different atom types, except for polymer-water oxygen interactions (see SI for details).
@@
[…]
@@
Figure 3: Arginine-water interactions. (a) Hydrogen bonds observed between backbone groups (${}^{-}$, NH+${}_{3}^{+}$) and the guanidinium (${}^{+}$) sidechain with water. (b) Hydrogen bond existence correlation functions for water-water, ${}^{-}$-water, ${}^{+}$-water, and NH+${}_{3}^{+}$-water. (c) Radial distribution function between OW and either CA (purple) or CZ (gold). (c, inset) Arrows denote trends observed with increasing arginine concentration. (d) A representative snapshot of hydrogen-bonding interactions involving arginine and water. Water molecules interacting with the ${}^{+}$ sidechain are highlighted in yellow, while those interacting with NH+${}_{3}^{+}$ and ${}^{-}$ are shaded in blue and purple, respectively.

### v2

@@ Abstract
Maintaining native protein structures in biological formulations poses a challenge, and is commonly addressed by strategic additive incorporation.1, 2, 3 Arginine stands out as a frequently employed additive in such formulations, spanning both therapeutic proteins4 and vaccines.5, 6 Arginine has been widely used as an aggregation suppressor, an agent for protein refolding, a cryoprotectant during lyophilization, and in protein purification. 7, 8, 9 Once hailed as a universal stabilizer, emerging studies paint a foggier picture of the effects of arginine. In some settings, the presence of arginine has accelerated the aggregation 10, 11, 12, denaturation13, 14, 15, and inactivation16 of certain proteins and viruses. Additional studies have found that arginine increases or decreases protein melting temperature depending on concentration 17, 18. Hence, the existing literature on arginine reveals a lack of a cohesive understanding of its multi-faceted effects on protein stability.
@@
We simulated a hydrophobic polymer in arginine solutions at different concentrations (Table S2). Replica exchange umbrella sampling60 (REUS) simulations were performed using GROMACS 2021.461, 62 with the PLUMED 2.8.0 63, 64 patch applied. The hydrophobic polymer was modeled as a linear coarse-grained chain with 26 monomers, where each monomer represents a $\mathrm{CH_{2}}$ unit with Lennard-Jones parameters $\sigma=0.373$ $\mathrm{nm}$ and $\epsilon=0.5856$ $\mathrm{kJ/mol}$38. Chloride (Cl-) counterions were added to achieve charge neutrality. The TIP4P/200565 model was used to describe water, and the CHARM22 force field was used for arginine and Cl- 66. Lorentz-Berthelot mixing rules67 were used to calculate non-bonded interactions between different atom types, except for polymer-water oxygen interactions (see SI for details).
@@
Fig. 3 describes hydrogen bonding interactions between arginine and water. Hydrogen bonds were calculated according to geometric criteria of a donor-acceptor distance of $\mathrm{r<=}$ 0.35 $\mathrm{nm}$ and an angle deviating less than 30o from 180o.88 Hydrogen bond existence correlation functions for water-water, guanidinium-water, carboxylate-water, and amine-water were estimated according to89, 90, 91
@@
Figure 3: Arginine-water interactions. (a) Hydrogen bonds observed between backbone groups (COO-, NH+${}_{3}^{+}$) and the guanidinium (Gdm+) sidechain with water. (b) Hydrogen bond existence correlation functions for water-water, COO--water, Gdm+-water, and NH+${}_{3}^{+}$-water. (c) Radial distribution function between OW and either CA (purple) or CZ (gold). (c, inset) Arrows denote trends observed with increasing arginine concentration. (d) A representative snapshot of hydrogen-bonding interactions involving arginine and water. Water molecules interacting with the Gdm+ sidechain are highlighted in yellow, while those interacting with NH+${}_{3}^{+}$ and COO- are shaded in blue and purple, respectively.